OpenStudy (anonymous):

Suppose f(x) = 23 + x and g(x) = x - 9 /3 Find the value of (g - f)(18).

OpenStudy (anonymous):

that would become (g-f)(18)=(18-9/3)-(23+18) all you do is replace the x with the value

OpenStudy (anonymous):

Okay thanks, I got the answer correct. How do we know that 18 is x?

OpenStudy (anonymous):

well if you look at (g-f)(18), this is the same thing as g(18)-f(18). g(x)=x-9/3 g(18)=18-9/3 f(x)=23+x f(18)=23+18 does that help?